OpenStudy (anonymous):

PLEASE HELP What is the solution to the equation log 3 x - log 3 2 = 2 ?

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

log(x)-log(2) = 2 log(x) = 2+log(2) x = 3^(2+log(2)) = 3^2*3^log(2) = 9*2 = 18

OpenStudy (zzr0ck3r):

I mean log base 3 when I write log

OpenStudy (anonymous):

THANK YOU SO MUCH! :)

OpenStudy (zzr0ck3r):

do you understand?

OpenStudy (zzr0ck3r):

another way 2 = log(x)-log(2) = log(x/2) so log(x/2) = 2 x/2 = 3^2 = 9 x=9*2=18